# TDD Without Tears: Towards Test Case Generation from Requirements through Deep Reinforcement Learning

### v1

@@ 3. PyTester
Aim. We introduce PyTester, a Text-to-Testcase generation approach that aims to automatically generate syntactically correct, executable, complete, and effective test cases based on a given natural language requirement. We formulate the Text-to-Testcase generation task as a Reinforcement Learning (RL) problem due to the following reasons:
@@
  • To increase models’ ability to generate alternative test cases. Traditionally, supervised finetuning approaches are trained to generate test cases that are textually and exactly matched with the ground truth test cases. Thus, these models can only generate test cases based on what the model had seen in the training data, limiting the model’s ability to generate test cases outside of the domain (i.e., out-of-distribution (OOD) problems (Shen et al., 2021)). In fact, the ground truth test cases may not be representative of all of the possible test cases. Thus, alternative test cases with different test inputs that still satisfy the given requirement should also be considered as valid.
@@
Figure 2. The formulation of Text-to-Testcase generation as a Reinforcement Learning problem.
PyTester leverages a Deep Reinforcement Learning (Deep RL) framework, which is a combination of deep learning and reinforcement learning principles. In the context of the Text-to-Testcase generation, the goal of RL is to enable the agent (i.e. the Text-to-Testcase generation model) and its policy (the brain of the agent) to learn by interacting with the environment (through trial and error) and receiving rewards (negative or positive) as feedback for performing actions (i.e. generating test cases). This formulation enables PyTester to (1) learn autonomously through interaction with their environment without ground truth test cases, allowing the models to generate alternative test cases and (2) learn how to generate syntactically correct, executable, complete, and effective test cases while being aligned with a given requirement through a carefully-crafted reward function rather than a basic loss function used in traditional finetuning models.
3.1. Our Reinforcement Learning Framework for Text-to-Testcase Generation
@@
Language Model: Given the natural language requirement $x$ and a testcase $y$, the Text-to-Testcase generation language model is defined as $P(y|x)=\prod_{i=0}^{m}p_{\theta}(y_{i}|x,y_{0:i-1})$
Dataset: We define a training dataset $\mathcal{D}$ of size $M$, which contains triplets of $<$requirement, code, test cases$>$ as $\mathcal{D}=\{(x^{i},c^{i},y^{i})\}_{i=1}^{i=M}$. The requirement $x$ consists of a signature of the Python function, the natural language requirement included in the docstring, the hidden code implementation (pass), the instruction prompt as a natural language comment “# check the correctness of the ‘...’ function”, and an incompleted assert statement “assert”. On the other hand, the ground truth test case $y$ consists of an assertion statement.
@@
State: A state $s_{t}=(x,\hat{y}_{0:t-1})$ is defined as a concatenation of the requirement $x$ and the test case tokens $\hat{y}_{0:t-1}$ generated so far.
@@
With the Markov Decision Process (MDP), our PyTester agent is learned as an episodic task. Generally, an episode is a sequence of interactions between the agent and environment starting from an initial state $s_{0}$ and ending at a terminal state where the agent will receive the reward and update the policy. At the beginning of each episode, the initial state will start from a requirement $x$ sampled from the Dataset $\mathcal{D}$, denoted as $s_{0}=x\sim\mathcal{D}_{x}$. Then, our policy $\pi_{\phi}$ takes a requirement $x$ as input to generate a test case $\hat{y}$ that is syntactically correct, executable, and complete. Technically, the policy $\pi_{\phi}$ will sequentially generate a sequence of tokens $\hat{y}_{i}$ for the test case $\hat{y}$ until reaching the end of the trajectory (i.e., after generating the EOS token or reaching the maximum generation length). Then, the generated test cases $\hat{y}$ are input into the interactive environment to receive a reward via a Reward Calculation function $R(x,\hat{y})$, which returns a cumulative reward point. The reward will be provided if the generated test cases are syntactically correct, executable against the ground truth code (i.e., aligned with the requirements), and complete (i.e., achieving high code coverage). On the other hand, the penalty will be provided if the generated test cases are syntactically incorrect and non-executable (e.g., runtime errors like incorrect method names and assertion errors like wrong test inputs). Finally, the policy $\pi_{\phi}$ is learned over time in order to maximize a reward function written as:
(1) $\mathcal{J}(\pi_{\phi})=\mathbb{E}_{x\sim\mathcal{D}_{x},\hat{y}\sim\pi_{\phi}(\cdot|x)}R(x,\hat{y})$
@@
Overview. Figure 3 presents an overview of our PyTester framework. Our framework consists of two steps: policy training and policy optimization. In the policy training step, we aim to build the agent (i.e., the Text-to-Testcase Generation model) where its policy interacts with the environment (i.e., the input dataset) and learns to make decisions (i.e., generate test cases) over time. However, currently, the agent is still suboptimal as it is trained to generate the same test cases as the ground truths in a supervised manner. Thus, the policy optimization step aims to further optimize the agent via a reward function through Deep Reinforcement Learning. To do so, given an environment (i.e., a triplet of $<$requirement, code, test cases$>$ in the training dataset), the agent will take requirement as input into the model in order to perform an action (i.e., generate test cases), where the test cases will be evaluated via a reward function. Our reward function is designed to incorporate three aspects of test case characteristics, namely syntax correctness, test executability, and test completeness. The agent is optimized with Proximal Policy Optimization (PPO) (Schulman et al., 2017) using our reward function, and stabilized with KL-divergence (Kullback and Leibler, 1951). Below, we describe the technical details of policy training, policy optimization, and our reward function.
@@
Figure 3. An overview of our Deep Reinforcement Learning framework for the Text-to-Testcase Generation task (called PyTester). ||||| | denotes a concatenation, $+$ denotes an addition, and -- denotes a subtraction.
3.2. Policy Training
@@
The ultimate goal of the policy training is to build a language model with some preliminary knowledge of the relationship between the requirement and its associated test cases. Therefore, we adopt the default CodeT5-large as our language model (Le et al., 2022; Raffel et al., 2020). To ensure that the model performs well on a target task (i.e., Text-to-Testcase generation), we finetune our language model with the training dataset. Given a training dataset $\mathcal{D}=\{(x^{i},c^{i},y^{i})\}_{i=1}^{i=M}$, we finetune the language model to learn the mapping between the requirement $x$ and its corresponding test cases $y$ (not the ground truth code $c$) in order to minimize the multi-class cross-entropy loss $\mathcal{L}_{\text{FT}}=-\mathbb{E}_{\begin{subarray}{c}(x,y)\sim\mathcal{D}\\ n\in\{0,...,N\}\end{subarray}}\text{log }p_{\theta}(y_{n}|x,y_{0:n-1})$, where $N$ is the maximum length of the test cases. Then, the loss is optimized using an Adam optimizer (Kingma and Ba, 2014) (the hyperparameters settings are provided in the Appendix).
3.3. Policy Optimization
@@
(2) $\mathcal{J}_{\text{surr}}(\pi_{\phi})=\text{min}$ $(rA^{\pi_{\phi_{\text{old}}}}(x,\hat{y}),\text{clip}(r,1-\epsilon,1+\epsilon)A^{\pi_{\phi_{\text{old}}}}(x,\hat{y})),$
(3) $\text{where}\quad r$ $=\frac{\pi_{\phi}(\hat{y}|x)}{\pi_{\phi_{\text{old}}}(\hat{y}|x)},$
in which $\epsilon$ is a small positive scalar indicating how much the policy can be updated and $\pi_{\phi_{\text{old}}}$ is the policy that collects the data. $A^{\pi}(x,\hat{y})$ is the advantage function of a policy $\pi$ given a state-action (i.e., input text requirement - output test cases) pair $(x,\hat{y})$, and is computed by the Generalized Advantage Estimator (GAE) (Schulman et al., 2015).
@@
(4) $R(x,\hat{y})=R_{\text{task}}-KL_{\text{penalty}}=R_{\text{task}}-\beta\text{log}\frac{\pi_{\phi}(\hat{y}|x)}{p_{\theta}(\hat{y}|x)},$
@@ 3.4. Reward Functions
(5) $R_{task}=R_{syntax}+R_{exec}+R_{cc}-P_{syntax}-P_{exec}-P_{assert}$
@@
(6) $R_{syntax}$ $=\mathds{1}_{syntax}r_{syntax},$
(7) $P_{syntax}$ $=(1-\mathds{1}_{syntax})p_{syntax},$
@@
(8) $R_{exec}$ $=\mathds{1}_{syntax}\mathds{1}_{exec}\mathds{1}_{assert}r_{exec},$
(9) $P_{exec}$ $=\mathds{1}_{syntax}(1-\mathds{1}_{exec})p_{exec},$
(10) $P_{assert}$ $=\mathds{1}_{syntax}\mathds{1}_{exec}(1-\mathds{1}_{assert})p_{assert}$
@@
(11) $R_{cov}=\mathds{1}_{syntax}\mathds{1}_{exec}\mathds{1}_{assert}r_{cov}$
4. Experimental Setup
@@
    Motivation. The reward function plays an important role in providing a positive/negative signal to the agent if the action satisfies the goal. There exist multiple variations of the reward functions (e.g., considering all types of feedback or considering only one type of feedback). Yet, little is known about which types of feedback are most important.
  3. (RQ3)
@@
    Motivation. Our PyTester approach may generate both correct and incorrect test cases. Thus, a manual analysis of the incorrectly generated test cases may offer opportunities for future researchers to consider improving the performance of our approach. Yet, little is known about what are the most common types of runtime errors that are found in the incorrectly generated test cases by PyTester.
4.1. Studied Dataset
The APPS benchmark dataset (Hendrycks et al., 2021) is used to evaluate our approach and compare it with the baseline approaches. The APPS benchmark dataset consists of 10,000 programming tasks with 131,777 associated test cases collected from Codewars, AtCoder, Kattis, and Codeforces. The dataset consists of three difficulty levels of the programming tasks, including the introductory level, the interview level, and the coding competition level. Each of the programming tasks consists of triplets of $<$a requirement, an associated code, and associated test cases$>$. We note that, in this study, only the APPS benchmark dataset is suitable for our Text-to-Testcase generation task since our PyTester model requires a large enough size of the dataset for policy training. Thus, other datasets like MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) are not considered in this experiment, as they are specifically designed for code generation, they are considered relatively small when compared to the APPs benchmark dataset, and are mainly designed for evaluating the zero-shot LLMs, not our Deep RL model. Similar to prior studies (Hendrycks et al., 2021), we use the same dataset splitting of 5,000 samples for training and 5,000 samples for testing.
4.2. Data Preprocessing
@@
Data Filtering. In addition, we found that not all samples in the dataset are of high quality. Common issues include missing test cases, missing code solutions, and non-executable test cases. Since the goal of PyTester is to generate test cases that are aligned with the requirement, our selection criteria are defined as the test cases that must exist and be executable against the code solution. After the data filtering step, our training dataset consists of 3,401 samples and our testing dataset consists of 3,259 samples.
4.3. Evaluation Metrics
Since the goal of PyTester is to generate syntactically correct, executable, complete, and effective test cases that are aligned with the requirement, our evaluation will focus on the following four aspects: syntax correctness, requirement alignment, code coverage, and mutation score.
@@ 4.3. Evaluation Metrics
  • Requirement Alignment measures the percentage of the generated test cases that are aligned with the requirement. Since the ground truth code is the correct implementation of the requirement, any test cases that can be successfully executed against the ground truth code are considered correct (i.e., correctly align with the given requirement and call the correct functions to be tested with the right test inputs and the expected outputs). Thus, we define that a test case is aligned with a given requirement if the test case is executable with the ground truth code. To do so, we execute the concatenation of the ground truth code with the generated test cases using the exec built-in function. [^7] https://docs.python.org/3/library/functions.html#exec
  • Code Coverage measures the completeness of the generated test cases for a given requirement. Ideally, in addition to aligning with the requirements, the test cases must be complete (i.e. executing all lines of the ground truth code solution). Thus, we measure the line-level code coverage of the generated test cases using the coverage.py Python library. [^8] https://coverage.readthedocs.io/en/7.2.7/
  • Mutation Score measures the effectiveness of the generated test cases. High-quality test cases must be able to detect bugs if the code is slightly changed. Mutation testing aims to introduce intentional faults (mutations) into the code and check if the existing test cases can detect these faults (mutations). Then, the mutation score is measured as a percentage based on the number of mutations that were killed (detected) versus the total number of mutations introduced. We use the MutPy Python library [^9] https://pypi.org/project/MutPy/ to measure a mutation score with a timeout of 10 seconds per sample.
@@
4.4. Baselines
Since the existing test case generation approaches (discussed in Section 2) are not designed for TDD (i.e. do not take a requirement as input), it becomes challenging to fairly compare our PyTester with the existing ones. Since prior studies demonstrated that Large Language Models (LLMs) could be a viable solution for various SE tasks, we consider the following four baseline models.

### v2

@@ 3 PyTester
Aim. We introduce PyTester, a Text-to-Testcase generation approach that aims to automatically generate syntactically correct, executable, complete, and effective test cases based on a given natural language description. We formulate the Text-to-Testcase generation task as a Reinforcement Learning (RL) problem due to the following reasons:
@@
  • To increase models’ ability to generate alternative test cases. Traditionally, supervised finetuning approaches are trained to generate test cases that are textually and exactly matched with the ground truth test cases. Thus, these models can only generate test cases based on what the model had seen in the training data, limiting the model’s ability to generate test cases outside of the domain (i.e., out-of-distribution (OOD) problems (Shen et al., 2021)). In fact, the ground truth test cases may not be representative of all of the possible test cases. Thus, alternative test cases with different test inputs that still satisfy the given description should also be considered as valid.
@@
Figure 2: The formulation of Text-to-Testcase generation as a Reinforcement Learning problem.
PyTester leverages a Deep Reinforcement Learning (Deep RL) framework, which is a combination of deep learning and reinforcement learning principles. In the context of the Text-to-Testcase generation, the goal of RL is to enable the agent (i.e. the Text-to-Testcase generation model) and its policy (the brain of the agent) to learn by interacting with the environment (through trial and error) and receiving rewards (negative or positive) as feedback for performing actions (i.e. generating test cases). This formulation enables PyTester to (1) learn autonomously through interaction with their environment without ground truth test cases, allowing the models to generate alternative test cases and (2) learn how to generate syntactically correct, executable, complete, and effective test cases while being aligned with a given description through a carefully-crafted reward function rather than a basic loss function used in traditional finetuning models.
3.1 Our Reinforcement Learning Framework for Text-to-Testcase Generation
@@
Language Model: Given the natural language description $x$ and a testcase $y$, the Text-to-Testcase generation language model is defined as $P(y|x)=\prod_{i=0}^{m}p_{\theta}(y_{i}|x,y_{0:i-1})$
Dataset: We define a training dataset $\mathcal{D}$ of size $M$, which contains triplets of $<$description, code, test cases$>$ as $\mathcal{D}=\{(x^{i},c^{i},y^{i})\}_{i=1}^{i=M}$. The description $x$ consists of a signature of the Python function, the natural language description included in the docstring, the hidden code implementation (pass), the instruction prompt as a natural language comment “# check the correctness of the ‘...’ function”, and an incompleted assert statement “assert”. On the other hand, the ground truth test case $y$ consists of an assertion statement.
@@
State: A state $s_{t}=(x,\hat{y}_{0:t-1})$ is defined as a concatenation of the textual description $x$ and the test case tokens $\hat{y}_{0:t-1}$ generated so far.
@@
With the Markov Decision Process (MDP), our PyTester agent is learned as an episodic task. Generally, an episode is a sequence of interactions between the agent and environment starting from an initial state $s_{0}$ and ending at a terminal state where the agent will receive the reward and update the policy. At the beginning of each episode, the initial state will start from a description $x$ sampled from the Dataset $\mathcal{D}$, denoted as $s_{0}=x\sim\mathcal{D}_{x}$. Then, our policy $\pi_{\phi}$ takes a description $x$ as input to generate a test case $\hat{y}$ that is syntactically correct, executable, and complete. Technically, the policy $\pi_{\phi}$ will sequentially generate a sequence of tokens $\hat{y}_{i}$ for the test case $\hat{y}$ until reaching the end of the trajectory (i.e., after generating the EOS token or reaching the maximum generation length). Then, the generated test cases $\hat{y}$ are input into the interactive environment to receive a reward via a Reward Calculation function $R(x,\hat{y})$, which returns a cumulative reward point. The reward will be provided if the generated test cases are syntactically correct, executable against the ground truth code (i.e., aligned with the descriptions), and complete (i.e., achieving high code coverage). On the other hand, the penalty will be provided if the generated test cases are syntactically incorrect and non-executable (e.g., runtime errors like incorrect method names and assertion errors like wrong test inputs). Finally, the policy $\pi_{\phi}$ is learned over time in order to maximize a reward function written as:
$\mathcal{J}(\pi_{\phi})=\mathbb{E}_{x\sim\mathcal{D}_{x},\hat{y}\sim\pi_{\phi}(\cdot|x)}R(x,\hat{y})$ (1)
@@
Overview. Figure 3 presents an overview of our PyTester framework. Our framework consists of two steps: policy training and policy optimization. In the policy training step, we aim to build the agent (i.e., the Text-to-Testcase Generation model) where its policy interacts with the environment (i.e., the input dataset) and learns to make decisions (i.e., generate test cases) over time. However, currently, the agent is still suboptimal as it is trained to generate the same test cases as the ground truths in a supervised manner. Thus, the policy optimization step aims to further optimize the agent via a reward function through Deep Reinforcement Learning. To do so, given an environment (i.e., a triplet of $<$description, code, test cases$>$ in the training dataset), the agent will take the description as input into the model in order to perform an action (i.e., generate test cases), where the test cases will be evaluated via a reward function. Our reward function is designed to incorporate three aspects of test case characteristics, namely syntax correctness, test executability, and test completeness. The agent is optimized with Proximal Policy Optimization (PPO) (Schulman et al., 2017) using our reward function, and stabilized with KL-divergence (Kullback and Leibler, 1951). Below, we describe the technical details of policy training, policy optimization, and our reward function.
@@
Figure 3: An overview of our Deep Reinforcement Learning framework for the Text-to-Testcase Generation task (called PyTester). ||||| | denotes a concatenation, $+$ denotes an addition, and -- denotes a subtraction.
3.2 Policy Training
@@
The ultimate goal of the policy training is to build a language model with some preliminary knowledge of the relationship between the description and its associated test cases. Therefore, we adopt the default CodeT5-large as our language model (Le et al., 2022; Raffel et al., 2020). To ensure that the model performs well on a target task (i.e., Text-to-Testcase generation), we finetune our language model with the training dataset. Given a training dataset $\mathcal{D}=\{(x^{i},c^{i},y^{i})\}_{i=1}^{i=M}$, we finetune the language model to learn the mapping between the description $x$ and its corresponding test cases $y$ (not the ground truth code $c$) in order to minimize the multi-class cross-entropy loss $\mathcal{L}_{\text{FT}}=-\mathbb{E}_{\begin{subarray}{c}(x,y)\sim\mathcal{D}\\ n\in\{0,...,N\}\end{subarray}}\text{log }p_{\theta}(y_{n}|x,y_{0:n-1})$, where $N$ is the maximum length of the test cases. Then, the loss is optimized using an Adam optimizer (Kingma and Ba, 2014) (the hyperparameters settings are provided in the Appendix).
3.3 Policy Optimization
@@
$\mathcal{J}_{\text{surr}}(\pi_{\phi})=\text{min}$ $(rA^{\pi_{\phi_{\text{old}}}}(x,\hat{y}),\text{clip}(r,1-\epsilon,1+\epsilon)A^{\pi_{\phi_{\text{old}}}}(x,\hat{y})),$ (2)
$\text{where}\quad r$ $=\frac{\pi_{\phi}(\hat{y}|x)}{\pi_{\phi_{\text{old}}}(\hat{y}|x)},$ (3)
in which $\epsilon$ is a small positive scalar indicating how much the policy can be updated and $\pi_{\phi_{\text{old}}}$ is the policy that collects the data. $A^{\pi}(x,\hat{y})$ is the advantage function of a policy $\pi$ given a state-action (i.e., input text description - output test cases) pair $(x,\hat{y})$, and is computed by the Generalized Advantage Estimator (GAE) (Schulman et al., 2015).
@@
$R(x,\hat{y})=R_{\text{task}}-KL_{\text{penalty}}=R_{\text{task}}-\beta\text{log}\frac{\pi_{\phi}(\hat{y}|x)}{p_{\theta}(\hat{y}|x)},$ (4)
@@ 3.4 Reward Functions
$R_{task}=R_{syntax}+R_{exec}+R_{cc}-P_{syntax}-P_{exec}-P_{assert}$ (5)
@@
$R_{syntax}$ $=\mathds{1}_{syntax}r_{syntax},$ (6)
$P_{syntax}$ $=(1-\mathds{1}_{syntax})p_{syntax},$ (7)
@@
$R_{exec}$ $=\mathds{1}_{syntax}\mathds{1}_{exec}\mathds{1}_{assert}r_{exec},$ (8)
$P_{exec}$ $=\mathds{1}_{syntax}(1-\mathds{1}_{exec})p_{exec},$ (9)
$P_{assert}$ $=\mathds{1}_{syntax}\mathds{1}_{exec}(1-\mathds{1}_{assert})p_{assert}$ (10)
@@
$R_{cov}=\mathds{1}_{syntax}\mathds{1}_{exec}\mathds{1}_{assert}r_{cov}$ (11)
4 Experimental Setup
@@
    Motivation. The reward function plays an important role in providing a positive/negative signal to the agent if the action satisfies the goal. There exist multiple variations of the reward functions (e.g., considering either one or multiple types of feedback). Yet, little is known about which types of feedback are most important.
  3. RQ3.
@@
    Motivation. Our PyTester approach may generate both correct and incorrect test cases. Thus, a manual analysis of the incorrectly generated test cases may offer opportunities for future researchers to consider improving the performance of our approach. Yet, little is known about what are the most common types of errors that are found in the incorrectly generated test cases by PyTester.
4.1 Studied Dataset
The APPS benchmark dataset (Hendrycks et al., 2021) is used to evaluate our approach and compare it with the baseline approaches. The APPS benchmark dataset consists of 10,000 programming tasks with 131,777 associated test cases collected from Codewars, AtCoder, Kattis, and Codeforces. The dataset consists of three difficulty levels of the programming tasks, including the introductory level, the interview level, and the coding competition level. Each of the programming tasks consists of triplets of $<$a description, an associated code, and associated test cases$>$. Similar to prior studies (Hendrycks et al., 2021), we use the same dataset splitting of 5,000 samples for training and 5,000 samples for testing.
margin:
Notably, in this study, only the APPS benchmark dataset is suitable for training our Text-to-Testcase generation task, as our PyTester model requires a large enough size of the dataset for policy training. Nevertheless, to broaden our evaluation, we provide additional analysis of PyTester in a zero-shot setting using MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) datasets. More details can be found in the discussion section (see Section 6).
4.2 Data Preprocessing
@@
Data Filtering. In addition, we found that not all samples in the dataset are of high quality. Common issues include missing test cases, missing code solutions, and non-executable test cases. Since the goal of PyTester is to generate test cases that are aligned with the natural language description, our selection criteria are defined as the test cases that must exist and be executable against the code solution. After the data filtering step, our training dataset consists of 3,401 samples and our testing dataset consists of 3,259 samples. During the model training step, we use 10% of the training dataset for validation, while the rest of the training dataset is used for model training.
4.3 Evaluation Metrics
Since the goal of PyTester is to generate syntactically correct, executable, complete, and effective test cases that are aligned with the description, our evaluation will focus on the following four aspects: syntax correctness, passing test case, code coverage, and mutation score.
@@ 4.3 Evaluation Metrics
  • Passing Test Case measures the percentage of the passing generated test cases to the ground truth code. Since the ground truth code is the correct implementation of the natural language description, any test cases that can be successfully executed against the ground truth code are considered passed (i.e., call the correct functions to be tested with the right test inputs and the expected outputs). Thus, we measure the passing test case by executing the concatenation of the ground truth code with the generated test cases using the exec built-in function. [^6] https://docs.python.org/3/library/functions.html#exec
  • Code Coverage measures the completeness of the generated test cases for a given natural language description. Ideally, in addition to the passing test case, the test cases must be complete (i.e. executing all lines of the ground truth code solution). Thus, we measure the line-level code coverage of the generated test cases using the coverage.py Python library. [^7] https://coverage.readthedocs.io/en/7.2.7/
  • Mutation Score measures the effectiveness of the generated test cases. High-quality test cases must be able to detect bugs if the code is slightly changed. Mutation testing aims to introduce intentional faults (mutations) into the code and check if the existing test cases can detect these faults (mutations). Then, the mutation score is measured as a percentage based on the number of mutations that were killed (detected) versus the total number of mutations introduced. margin: We use the MutPy Python library [^8] https://pypi.org/project/MutPy/ to measure a mutation score with a timeout of 30 seconds per sample. We set the configuration to default with PyTest [^9] https://docs.pytest.org/en/latest/ as a test runner. The mutation operators are listed in Table 1.
Table 1: The operators used in our mutation testing.
Mutation Testing Operator
AOD - arithmetic operator deletion
AOR - arithmetic operator replacement
ASR - assignment operator replacement
BCR - break continue replacement
COD - conditional operator deletion
COI - conditional operator insertion
CRP - constant replacement
DDL - decorator deletion
EHD - exception handler deletion
EXS - exception swallowing
IHD - hiding variable deletion
IOD - overriding method deletion
IOP - overridden method calling position change
LCR - logical connector replacement
LOD - logical operator deletion
LOR - logical operator replacement
ROR - relational operator replacement
SCD - super calling deletion
SCI - super calling insert
SIR - slice index remove
@@
4.4 Evaluation Setting
Since the number of test cases generated from each model varies, we provide 3 evaluation settings: Raw, n-test, and Filtered by Syntax Correctness, to ensure a fair comparison.
  • Raw evaluates the generated test cases directly from the model without any post-processing. In this setting, the test cases are assessed as a whole set. If any single test case fails, the entire output is considered incorrect.
  • n-test evaluates the generated test cases by limiting the size to n test cases. We use n=3 (3-test) and n=5 (5-test).
  • Filtered by Syntax Correctness evaluates the generated test cases after filtering out the incorrect syntax of the test cases. First, we split multiple assertions into individual tests, then exclude the syntactically incorrect test cases before evaluation.
4.5 Baselines
Since the existing test case generation approaches (discussed in Section 2) are not designed for TDD (i.e. do not take a description as input), it becomes challenging to fairly compare our PyTester with the existing ones. Since prior studies demonstrated that Large Language Models (LLMs) could be a viable solution for various SE tasks, we consider the following five baseline models. Note that if available, we use beam search with a size of 5 as the decoding method. Otherwise, we conduct the experiments five times per data point and report the mean and standard deviation.